# Pre-trained Transformer Uncovers Meaningful Patterns in Human Mobility Data

### v1

@@ Abstract.
1. Introduction
Transformers pre-trained on internet-scale text have significantly advanced the field of Natural Language Processing (NLP) over the past few years (Devlin et al., 2018; Zhang et al., 2019; Liu et al., 2019; Lan et al., 2019; Sanh et al., 2019; Clark et al., 2020; He et al., 2020; Radford et al., 2018, 2019; Yang et al., 2019; Brown et al., 2020; Workshop et al., 2022; Zhang et al., 2022; Touvron et al., 2023). Whether pre-trained to predict the next word in a sentence (Radford et al., 2018, 2019; Yang et al., 2019; Brown et al., 2020; Workshop et al., 2022; Zhang et al., 2022; Touvron et al., 2023) or unmask randomly masked text (Devlin et al., 2018; Zhang et al., 2019; Liu et al., 2019; Lan et al., 2019; Sanh et al., 2019; Clark et al., 2020; He et al., 2020), these models learn rich word embeddings that excel at a wide range of language tasks (Devlin et al., 2018; Radford et al., 2018).
Motivated by their success in NLP and recognizing the inherent parallels between human mobility data and natural language, the geospatial artificial intelligence (GeoAI) community has recently expressed keen interest in pre-trained transformers (Musleh et al., 2022; Musleh, 2022; Musleh and Mokbel, 2023; Najjar, 2023; Xue et al., 2022; Xue and Salim, 2023b; Zhang et al., 2023; Wang et al., 2023; Liang et al., 2023). Despite encountering challenges, such as limited pre-training data, these studies have demonstrated the potential of such models for human mobility understanding.
While this recent development marks a significant step forward in the field. It raises questions about whether transformers pre-trained on unlabeled human mobility data (Musleh et al., 2022; Musleh, 2022; Musleh and Mokbel, 2023; Najjar, 2023) truly capture meaningful mobility patterns, akin to how word embeddings capture structure and semantics in natural language.
In this paper, we attempt to answer this question by thoroughly evaluating embeddings learned by a bidirectional transformer pre-trained via random masking (Devlin et al., 2018) on real-world, country-scale, unlabeled human mobility data to encapsulate a wide array of concepts directly and indirectly related to human mobility.
@@ 1. Introduction
The remainder of this paper is organized into four key sections. In Section 2, we provide an overview of previous works that utilize pre-trained transformers for human mobility understanding. Section 3 outlines the materials and methods employed in our study, offering insights into data, transformer architecture, pre-training and adaptation approaches used. Section 4 presents the evaluation results of our pre-trained embeddings. Finally, in Section 5, we summarize the paper, draw conclusions, and engage in a discussion on our findings and avenues for future research.
@@
| Xue et al (Xue et al., 2022; Xue and Salim, 2023b) | Language | POI visitor forecasting |
| Wang et al (Wang et al., 2023) | Next-location prediction | |
| Zhang et al (Zhang et al., 2023) | Anomaly detection | |
| Liang et al (Liang et al., 2023) | Public event prediction | |
| Musleh et al (Musleh et al., 2022) | Trajectory | Task agnostic |
| Musleh (Musleh, 2022) | ||
| Musleh & Mokbel (Musleh and Mokbel, 2023) | ||
| Najjar (Najjar, 2023) |
2. Previous Works
@@
Previous works that utilize pre-trained transformers for human mobility understanding can be broadly categorized into two groups: studies utilizing transformers pre-trained on language, i.e., large language models (LLMs) (Xue et al., 2022; Xue and Salim, 2023b; Zhang et al., 2023; Wang et al., 2023; Liang et al., 2023), and those using transformers pre-trained from scratch on trajectory data (Musleh et al., 2022; Musleh, 2022; Musleh and Mokbel, 2023; Najjar, 2023). See Table 1 for a summary.
@@
In the first group, human mobility data is translated into natural language prompts which are employed to fine-tune an LLM on a target downstream task. Xue et al were the first to showcase the potential of this approach for human mobility understanding. In their work (Xue et al., 2022; Xue and Salim, 2023b), they empirically demonstrated that, for the task of Point-of-interest (POI) visitor forecasting, fine-tuned LLMs perform comparably to specialized models trained exclusively on numerical data.
@@
Other researchers followed suit, reporting similar results on different human mobility understanding tasks, such as next-location prediction (Wang et al., 2023), trajectory anomaly detection (Zhang et al., 2023) and public event prediction (Liang et al., 2023).
@@
The second group, on the other hand, is a part of an emergent effort within the GeoAI community aimed at developing Geo Foundation Models (Mai et al., 2022, 2023a; Xue and Salim, 2023a; Rao et al., 2023; Mai et al., 2023b) for trajectory intelligence. These models are task-agnostic, pre-trained from scratch on extensive trajectory data with the goal of developing a profound understanding of the dynamics of human mobility.
@@
In the work by Musleh et al (Musleh et al., 2022), a vision for a unified framework for trajectory analysis is proposed. The authors advocate for a model pre-trained once on unlabeled trajectory data, and minimally fine-tuned for a wide range of downstream tasks. This vision is realized in the framework named TrajBERT, as demonstrated by Musleh in (Musleh, 2022), where a BERT-like transformer, pre-trained on taxi-trip data, learns to impute trajectories. Subsequently, Musleh and Mokbel reported similar results on a larger dataset in (Musleh and Mokbel, 2023). Later, Najjar (Najjar, 2023) demonstrated that a BERT-like transformer, pre-trained through random masking on user check-in data, not only acquires trajectory unmasking skills but also seems to learn higher order tasks, such as trajectory-user linking (Najjar and Mede, 2022).
@@
While our work resembles (Musleh et al., 2022; Musleh, 2022; Musleh and Mokbel, 2023; Najjar, 2023) in terms of training objective, transformer architecture and data modality, it is to the best of our knowledge the first to present extensive empirical evidence suggesting that a transformer pre-trained on unlabeled human mobility data learns embeddings capable, through fine-tuning, of developing a deep understanding of high-level geospatial attributes, such as population and land cover, of the target study area.
3. Materials and Methods
3.1. Data
@@
We utilize a proprietary dataset of location data, collected and maintained by Rakuten111https://www.rakuten.com/. The dataset contains over 53 billion GPS data points generated by more than 6 million anonymous users active through out Japan between July 2022 and June 2023. See Figure 2 for a visualization of the dataset’s spatial coverage. We processed this data into approximately 17 million trajectories, which, to the best of our knowledge, is significantly larger than any other public dataset of its kind (Cho et al., 2011; Yang et al., 2016; Liu, Yong, 2022; ACM, 2017). We used this dataset to pre-train, validate, and evaluate our model as explained in the following. See Appendix A.1 for details on data processing.
3.2. Transformer
@@
We adopt BERT (Devlin et al., 2018) as the base architecture for our pre-trained transformer. BERT, which stands for Bidirectional Encoder Representations from Transformers, is a transformer-based model released by Google in 2018. It has revolutionized the field of NLP by introducing a bidirectional approach to language modeling. Unlike unidirectional models, BERT considers context from both directions, enabling it to grasp the meaning of words based on their surrounding context. This bidirectional understanding enhances the model’s ability to comprehend nuances, leading to improved performance in various NLP tasks.
Similar to previous works (Musleh et al., 2022; Musleh, 2022; Musleh and Mokbel, 2023; Najjar, 2023), we adapt BERT to handle trajectory data by incorporating a spatial-tokenization step which is necessary since transformers are designed to process natural language rather than location data (Najjar, 2023). See Appendix A.2 for implementation details.
3.3. Pre-training
@@
We leverage Masked Trajectory Modeling (MTM) as a self-supervised pre-training task, a technique previously employed for the same purpose in (Najjar, 2023). A related pre-training task, trajectory imputation, was utilized by Musleh et al in (Musleh et al., 2022; Musleh, 2022; Musleh and Mokbel, 2023). Both approaches have demonstrated the potential of random masking for learning from unlabeled trajectory data.
@@
Drawing parallels with Masked Language Modeling (MLM) in NLP (Devlin et al., 2018), MTM involves randomly masking a portion of a trajectory, prompting the model to predict the missing points based on their surrounding context. See Figure 1 for an illustration.
@@
The process of learning to unmask millions of trajectories presumably forces the model to learn embeddings that capture the contextual relationships between the hashes constituting the target study area. This rationale underscores our choice of employing MTM as a pre-training task. See Appendix A.3 for implementation details.
3.4. Adaptation
@@
To this end, we utilize three well-known adaptation approaches, namely fine-tuning, few-shot learning and zero-shot learning. See Appendix A.3 for implementation details.
4. Results
We present the obtained adaptation results organized into two groups based on the model’s input whether it is a region or a trajectory. Refer to Appendix A.4 for details on evaluation metrics, and pre-training/adaptation implementation.
4.1. Regions
Learning embeddings to represent geographic regions is an active area of GeoAI research (Spruyt, 2018; Du et al., 2018; Jenkins et al., 2019; Wang et al., 2020; Xiang, 2020; Woźniak and Szymański, 2021). In a similar vein, we demonstrate that embeddings pre-trained on randomly masked sequences of (geo) hashes can effectively encode various attributes of the regions these hashes represent. Specifically, we examine the following: geographic location, population count, administrative divisions, and land cover. See Appendix A.3 for details on the datasets used to adapt different models.
@@ 4.1. Regions
The obtained results, summarized in Table 2, indicate that pre-training (compared to initialization with random weights) reduces inference error by 0.476 and 0.432 degree in latitude and longitude, respectively. Which translates to 71 km in terms of Haversine distance. This reduction is significant, especially considering that we are using hexagons with an area of 0.74 km2 each. It is worth noting that we used relative coordinates, meaning coordinates were calculated with respect to a reference point, which is the central point of the study area.
@@
Population count refers to the total number of individuals within a specific geographic area. In the following, we assess our model’s ability to infer population at the hexagon level. To this end, we utilize WorldPop (WorldPop and Center for International Earth Science Information Network , Columbia University(2018), CIESIN) data aggregated at a 1 km resolution as a ground truth. Since this data is not directly measured but rather collected using remote sensing methods, we frame the target task as a 10-class (deciles) classification rather than direct population count regression.
While the obtained results (refer to Table 3) do not suggest any of the models perform well at the task at hand, they clearly indicate that pre-training is beneficial, with overall gain of 0.115 in F-Score (33% performance gain). Furthermore, a map of the inferred population (see Figure 4) illustrates the model’s ability to preserve (to a certain degree) the spatial distribution of population over the study area.
| Model | Accuracy | Precision | Recall | F-Score |
@@
Administrative divisions
Japan, for example, consists of eight main regions divided into 47 prefectures (see Figure 5). Municipalities within each prefecture are assigned one of three broad categories: Shi (city), Machi (town), or Mura (village) (Wikipedia contributors, 2024).
Through a series of experiments, we evaluate how well our pre-trained embeddings can encapsulate attributes such as prefecture, region, and municipality. In all three cases, the problem is framed as hexagon classification with the objective of assigning a given hexagon its prefecture, region and municipality classification, respectively. We used GADM dataset (Areas, Global Administrative, 2024) as a ground truth.
Results are summarized in Tables 4, 5, and 6. The impact of pre-training is clearly noticeable in all three scenarios, with an overall performance boost ranging between 5% and 12% across the three classification tasks. Refer to Figure 6 for a subjective evaluation of the best model performance at municipality classification.
| Model | Accuracy | Precision | Recall | F-Score |
| TransformerSmall | 0.8 | 0.866 | 0.862 | 0.863 |
| TransformerMedium | 0.79 | 0.858 | 0.857 | 0.856 |
| TransformerLarge | 0.791 | 0.861 | 0.859 | 0.857 |
| BERTRandomly-initialized | 0.79 | 0.86 | 0.857 | 0.857 |
| BERTPre-trained (Zero-shot) | 0.024 | 0.003 | 0.017 | 0.003 |
| BERTPre-trained (Few-shots) | 0.021 | 0.034 | 0.185 | 0.058 |
| BERTPre-trained (Fine-tuned) | 0.958 | 0.969 | 0.968 | 0.968 |
| Model | Accuracy | Precision | Recall | F-Score |
| TransformerSmall | 0.932 | 0.939 | 0.938 | 0.938 |
| TransformerMedium | 0.933 | 0.94 | 0.937 | 0.938 |
| TransformerLarge | 0.934 | 0.942 | 0.936 | 0.938 |
| BERTRandomly-initialized | 0.934 | 0.942 | 0.94 | 0.941 |
| BERTPre-trained (Zero-shot) | 0.123 | 0.03 | 0.129 | 0.037 |
| BERTPre-trained (Few-shots) | 0.125 | 0.046 | 0.178 | 0.054 |
| BERTPre-trained (Fine-tuned) | 0.994 | 0.994 | 0.994 | 0.994 |
| Model | Accuracy | Precision | Recall | F-Score |
@@
We evaluate the ability of our model to infer the distribution of two land cover classes, namely built-up areas and tree cover. We utilize a global land-cover land-use dataset sampled at a 10 m resolution (Karra et al., 2021). We labeled each hexagon by its ratio of the target land cover and framed the problem as regression. The results are summarized in Tables 7 and 8.
@@
In both cases, pre-training significantly boosts performance, with an overall gain of 0.147 to 0.158 in R2 score which translates to 26% to 38% improvement in variance explanation. For a subjective evaluation we mapped in Figures 7 and 8 the inferred distribution of the respective land cover classes along side ground-truth data. The model’s ability to preserve the spatial distribution is clearly illustrated in both examples.
@@ 4.2. Trajectories
The obtained results (refer to Table 9) show significant performance improvement from pre-training, with an R2 score boost of 0.715. This task builds upon prefecture classification, so the performance gain from pre-training aligns with expectations. However, the extent of this improvement is intriguing. We attribute it to BERT’s ability to leverage context, which is lacking when working with regions, e.g., prefecture classification.
@@
In this experiment, we evaluate the effectiveness of our pre-trained embeddings in encapsulating trip length. Our findings, summarized in Table 10, demonstrate a notable gain in R2 of over 0.7 (or 8.5 km reduction in inference error) as a result of pre-training. We attribute this performance boost to the pre-training’s ability to capture spatial relationships among the hexagons constituting the target study area, as evidenced in the geographic location experiment presented earlier.
@@
We are interested in two travel patterns: cross-prefecture and cross-region travel. Travel, like other offline activities, has been the focus of previous studies (Najjar and Mede, 2023). To ensure that we exclude commuting, we define travel as crossing prefecture/region borders during weekends and/or national holidays. To this end, we frame the problem as a simple binary classification task, aiming to infer whether the user is a traveler based on their activity over a month.
The obtained results, summarized in Table 11, not only indicate that our embeddings effectively encapsulate both travel patterns but also demonstrate that pre-training significantly enhances the model’s performance by up to 0.26 in F-Score. This finding adds to the evidence that random masking applied to trajectory data is highly effective in learning the administrative divisions of the target study area.
@@
Next, we evaluate our model’s capacity to infer the user’s exposure to greenery, as depicted in Table 12. Results indicate that pre-training has improved the model’s performance by over 40% in terms of R2 score.

### v2

@@ Abstract
Index Terms:
Geospatial artificial intelligence; geo foundation models; pre-trained transformersI Introduction
Transformers pre-trained on internet-scale text have significantly advanced the field of Natural Language Processing (NLP) over the past few years [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Whether pre-trained to predict the next word in a sentence [8, 9, 10, 11, 12, 13, 14] or unmask randomly masked text [1, 2, 3, 4, 5, 6, 7], these models learn rich word embeddings that excel at a wide range of language tasks [1, 8].
Motivated by their success in NLP and recognizing the inherent parallels between human mobility data and natural language, the geospatial artificial intelligence (GeoAI) community has recently expressed keen interest in pre-trained transformers [15, 16, 17, 18, 19, 20, 21, 22, 23]. Despite encountering challenges, such as limited pre-training data, these studies have demonstrated the potential of such models for human mobility understanding.
While this recent development marks a significant step forward in the field. It raises questions about whether transformers pre-trained on unlabeled human mobility data [15, 16, 17, 18] truly capture meaningful mobility patterns, akin to how word embeddings capture structure and semantics in natural language.
In this paper, we attempt to answer this question by thoroughly evaluating embeddings learned by a bidirectional transformer pre-trained via random masking [1] on real-world, country-scale, unlabeled human mobility data to encapsulate a wide array of concepts directly and indirectly related to human mobility.
@@
The remainder of this paper is organized into four key sections. In Section II, we provide an overview of previous works that utilize pre-trained transformers for human mobility understanding. Section III outlines the materials and methods employed in our study, offering insights into data, transformer architecture, pre-training and adaptation approaches used. Section IV presents the evaluation results of our pre-trained embeddings. Finally, in Section V, we summarize the paper, draw conclusions, and engage in a discussion on our findings and avenues for future research.
@@
| Xue et al [19, 20] | Language | POI visitor forecasting |
| Wang et al [22] | Next-location prediction | |
| Zhang et al [21] | Anomaly detection | |
| Liang et al [23] | Public event prediction | |
| Musleh et al [15] | Trajectory | Task agnostic |
| Musleh [16] | ||
| Musleh & Mokbel [17] | ||
| Najjar [18] |
II Previous Works
@@
Previous works that utilize pre-trained transformers for human mobility understanding can be broadly categorized into two groups: studies utilizing transformers pre-trained on language, i.e., large language models (LLMs) [19, 20, 21, 22, 23], and those using transformers pre-trained from scratch on trajectory data [15, 16, 17, 18]. See Table I for a summary.
@@
In the first group, human mobility data is translated into natural language prompts which are employed to fine-tune an LLM on a target downstream task. Xue et al were the first to showcase the potential of this approach for human mobility understanding. In their work [19, 20], they empirically demonstrated that, for the task of Point-of-interest (POI) visitor forecasting, fine-tuned LLMs perform comparably to specialized models trained exclusively on numerical data.
@@
Other researchers followed suit, reporting similar results on different human mobility understanding tasks, such as next-location prediction [22], trajectory anomaly detection [21] and public event prediction [23].
@@
The second group, on the other hand, is a part of an emergent effort within the GeoAI community aimed at developing Geo Foundation Models [24, 25, 26, 27, 28] for trajectory intelligence. These models are task-agnostic, pre-trained from scratch on extensive trajectory data with the goal of developing a profound understanding of the dynamics of human mobility.
@@
In the work by Musleh et al [15], a vision for a unified framework for trajectory analysis is proposed. The authors advocate for a model pre-trained once on unlabeled trajectory data, and minimally fine-tuned for a wide range of downstream tasks. This vision is realized in the framework named TrajBERT, as demonstrated by Musleh in [16], where a BERT-like transformer, pre-trained on taxi-trip data, learns to impute trajectories. Subsequently, Musleh and Mokbel reported similar results on a larger dataset in [17]. Later, Najjar [18] demonstrated that a BERT-like transformer, pre-trained through random masking on user check-in data, not only acquires trajectory unmasking skills but also seems to learn higher order tasks, such as trajectory-user linking [29].
@@
While our work resembles [15, 16, 17, 18] in terms of training objective, transformer architecture and data modality, it is to the best of our knowledge the first to present extensive empirical evidence suggesting that a transformer pre-trained on unlabeled human mobility data learns embeddings capable, through fine-tuning, of developing a deep understanding of high-level geospatial attributes, such as population and land cover, of the target study area.
III Materials and Methods
III-A Data
@@
We utilize a proprietary dataset of location data, collected and maintained by Rakuten111https://www.rakuten.com/. The dataset contains over 53 billion GPS data points generated by more than 6 million anonymous users active through out Japan between July 2022 and June 2023. See Figure 2 for a visualization of the dataset’s spatial coverage. We processed this data into approximately 17 million trajectories, which, to the best of our knowledge, is significantly larger than any other public dataset of its kind [30, 31, 32, 33]. We used this dataset to pre-train, validate, and evaluate our model as explained in the following. See Appendix A-A for details on data processing.
III-B Transformer
@@
We adopt BERT [1] as the base architecture for our pre-trained transformer. BERT, which stands for Bidirectional Encoder Representations from Transformers, is a transformer-based model released by Google in 2018. It has revolutionized the field of NLP by introducing a bidirectional approach to language modeling. Unlike unidirectional models, BERT considers context from both directions, enabling it to grasp the meaning of words based on their surrounding context. This bidirectional understanding enhances the model’s ability to comprehend nuances, leading to improved performance in various NLP tasks.
III-C Pre-training
@@
We leverage Masked Trajectory Modeling (MTM) as a self-supervised pre-training task, a technique previously employed for the same purpose in [18]. A related pre-training task, trajectory imputation, was utilized by Musleh et al in [15, 16, 17]. Both approaches have demonstrated the potential of random masking for learning from unlabeled trajectory data.
@@
Drawing parallels with Masked Language Modeling (MLM) in NLP [1], MTM involves randomly masking a portion of a trajectory, prompting the model to predict the missing points based on their surrounding context. See Figure 1 for an illustration.
@@
The process of learning to unmask millions of trajectories presumably forces the model to learn embeddings that capture the contextual relationships between the hashes constituting the target study area. This rationale underscores our choice of employing MTM as a pre-training task. See Appendix A-C for implementation details.
III-D Adaptation
@@
To this end, we utilize three well-known adaptation approaches, namely fine-tuning, few-shot learning and zero-shot learning. See Appendix A-C for implementation details.
IV Results
We present the obtained adaptation results organized into two groups based on the model’s input whether it is a region or a trajectory. Refer to Appendix A-D for details on evaluation metrics, and pre-training/adaptation implementation.
IV-A Regions
Learning embeddings to represent geographic regions is an active area of GeoAI research [34, 35, 36, 37, 38, 39]. In a similar vein, we demonstrate that embeddings pre-trained on randomly masked sequences of (geo) hashes can effectively encode various attributes of the regions these hashes represent. Specifically, we examine the following: geographic location, population count, administrative divisions, and land cover. See Appendix A-C for details on the datasets used to adapt different models.
@@ IV-A Regions
The obtained results, summarized in Table II, indicate that pre-training (compared to initialization with random weights) reduces inference error by 0.476 and 0.432 degree in latitude and longitude, respectively. Which translates to 71 km in terms of Haversine distance. This reduction is significant, especially considering that we are using hexagons with an area of 0.74 km2 each. It is worth noting that we used relative coordinates, meaning coordinates were calculated with respect to a reference point, which is the central point of the study area.
@@
Population count refers to the total number of individuals within a specific geographic area. In the following, we assess our model’s ability to infer population at the hexagon level. To this end, we utilize WorldPop [40] data aggregated at a 1 km resolution as a ground truth. Since this data is not directly measured but rather collected using remote sensing methods, we frame the target task as a 10-class (deciles) classification rather than direct population count regression.
While the obtained results (refer to Table V) do not suggest any of the models perform well at the task at hand, they clearly indicate that pre-training is beneficial, with overall gain of 0.115 in F-Score (33% performance gain). Furthermore, a map of the inferred population (see Figure 4) illustrates the model’s ability to preserve (to a certain degree) the spatial distribution of population over the study area.
Administrative divisions
Japan, for example, consists of eight main regions divided into 47 prefectures. Municipalities within each prefecture are assigned one of three broad categories: Shi (city), Machi (town), or Mura (village) [41].
Through a series of experiments, we evaluate how well our pre-trained embeddings can encapsulate attributes such as prefecture, region, and municipality. In all three cases, the problem is framed as hexagon classification with the objective of assigning a given hexagon its prefecture, region and municipality classification, respectively. We used GADM dataset [42] as a ground truth.
Results are summarized in Tables III, IV, and VI. The impact of pre-training is clearly noticeable in all three scenarios, with an overall performance boost ranging between 5% and 12% across the three classification tasks. Refer to Figure 5 for a subjective evaluation of the best model performance at municipality classification.
| Model | Acc. | P | R | F |
| TransformerSmall | 0.8 | 0.866 | 0.862 | 0.863 |
| TransformerMedium | 0.79 | 0.858 | 0.857 | 0.856 |
| TransformerLarge | 0.791 | 0.861 | 0.859 | 0.857 |
| BERTRandomly-initialized | 0.79 | 0.86 | 0.857 | 0.857 |
| BERTPre-trained (Zero-shot) | 0.024 | 0.003 | 0.017 | 0.003 |
| BERTPre-trained (Few-shots) | 0.021 | 0.034 | 0.185 | 0.058 |
| BERTPre-trained (Fine-tuned) | 0.958 | 0.969 | 0.968 | 0.968 |
| Model | Acc. | P | R | F |
| TransformerSmall | 0.932 | 0.939 | 0.938 | 0.938 |
| TransformerMedium | 0.933 | 0.94 | 0.937 | 0.938 |
| TransformerLarge | 0.934 | 0.942 | 0.936 | 0.938 |
| BERTRandomly-initialized | 0.934 | 0.942 | 0.94 | 0.941 |
| BERTPre-trained (Zero-shot) | 0.123 | 0.03 | 0.129 | 0.037 |
| BERTPre-trained (Few-shots) | 0.125 | 0.046 | 0.178 | 0.054 |
| BERTPre-trained (Fine-tuned) | 0.994 | 0.994 | 0.994 | 0.994 |
| Model | Acc. | P | R | F |
@@
| Model | Acc. | P | R | F |
@@
We evaluate the ability of our model to infer the distribution of two land cover classes, namely built-up areas and tree cover. We utilize a global land-cover land-use dataset sampled at a 10 m resolution [43]. We labeled each hexagon by its ratio of the target land cover and framed the problem as regression. The results are summarized in Tables VII and VIII.
@@
In both cases, pre-training significantly boosts performance, with an overall gain of 0.147 to 0.158 in R2 score which translates to 26% to 38% improvement in variance explanation. For a subjective evaluation we mapped in Figures 6 and 7 the inferred distribution of the respective land cover classes along side ground-truth data. The model’s ability to preserve the spatial distribution is clearly illustrated in both examples.
@@ IV-B Trajectories
The obtained results (refer to Table IX) show significant performance improvement from pre-training, with an R2 score boost of 0.715. This task builds upon prefecture classification, so the performance gain from pre-training aligns with expectations. However, the extent of this improvement is intriguing. We attribute it to BERT’s ability to leverage context, which is lacking when working with regions, e.g., prefecture classification.
@@
In this experiment, we evaluate the effectiveness of our pre-trained embeddings in encapsulating trip length. Our findings, summarized in Table X, demonstrate a notable gain in R2 of over 0.7 (or 8.5 km reduction in inference error) as a result of pre-training. We attribute this performance boost to the pre-training’s ability to capture spatial relationships among the hexagons constituting the target study area, as evidenced in the geographic location experiment presented earlier.
@@
We are interested in two travel patterns: cross-prefecture and cross-region travel. Travel, like other offline activities, has been the focus of previous studies [44]. To ensure that we exclude commuting, we define travel as crossing prefecture/region borders during weekends and/or national holidays. To this end, we frame the problem as a simple binary classification task, aiming to infer whether the user is a traveler based on their activity over a month.
The obtained results, summarized in Table XI, not only indicate that our embeddings effectively encapsulate both travel patterns but also demonstrate that pre-training significantly enhances the model’s performance by up to 0.26 in F-Score. This finding adds to the evidence that random masking applied to trajectory data is highly effective in learning the administrative divisions of the target study area.
@@
Next, we evaluate our model’s capacity to infer the user’s exposure to greenery, as depicted in Table XII. Results indicate that pre-training has improved the model’s performance by over 40% in terms of R2 score.